OpenStudy (anonymous):

Find i)The Vertex ii) Equation of Line Symetry y= (x-p)^2 +q

OpenStudy (goformit100):

@ParthKohli

OpenStudy (anonymous):

i, when x=p, the y will have the minimum value q. so the coordinate of the vertex is (p,q) ii, y-axis symmetry y=(p-x)^2+q x-axis symmetry y=-(x-p)^2-q

OpenStudy (anonymous):

sorry, something wrong. the second one, ii, y-axis symmetry should be y=(x+p)^2+q

OpenStudy (callisto):

Vertex is at (p,q) Axis of symmetry is x=p